# EMORL: Ensemble Multi-Objective Reinforcement Learning for Efficient and Flexible LLM Fine-Tuning

### v2

@@ Abstract
Recent advances in reinforcement learning (RL) for large language model (LLM) fine-tuning show promise in addressing multi-objective tasks but still face significant challenges, including complex objective balancing, low training efficiency, poor scalability, and limited explainability. Leveraging ensemble learning principles, we introduce an Ensemble Multi-Objective RL (EMORL) framework that fine-tunes multiple models with individual objectives while optimizing their aggregation after the training to improve efficiency and flexibility. Our method is the first to aggregate the last hidden states of individual models, incorporating contextual information from multiple objectives. This approach is supported by a hierarchical grid search algorithm that identifies optimal weighted combinations. We evaluate EMORL on counselor reflection generation tasks, using text-scoring LLMs to evaluate the generations and provide rewards during RL fine-tuning. Through comprehensive experiments on the PAIR and Psych8k datasets, we demonstrate the advantages of EMORL against existing baselines: significantly lower and more stable training consumption ($17,529\pm 1,650$ data points and $6,573\pm 147.43$ seconds), improved scalability and explainability, and comparable performance across multiple objectives.
@@
Lingxiao Kong1, Cong Yang2, Susanne Neufang3, Oya Deniz Beyan1,3, Zeyd Boukhers1,3 1Fraunhofer Institute for Applied Information Technology FIT, Germany 2Soochow University, China 3University Hospital of Cologne, Germany
@@ 1 Introduction
Multi-objective optimization for large language models (LLM) is a crucial research direction for natural language processing (NLP) tasks that need to fulfill diverse and often competing requirements (Vaswani, 2017; Qin et al., 2024; Kumar, 2024). Such tasks require models to simultaneously optimize for multiple evaluation criteria, balancing trade-offs between different objectives. This paper focuses on the counselor reflection generation task, where LLM reflects on users’ prompts with empathy, building the first step of therapeutic communication (O’neil et al., 2023). As exemplified in Figure 1, this task requires reflective responses that optimize multiple key objectives, such as reflection, empathy, and fluency in high-quality counseling interactions. We employ reinforcement learning (RL) to fine-tune pre-trained LLMs for these specific objectives (Lin, 2024; Parthasarathy et al., 2024). Conventional RL fine-tuning approaches combine multiple objectives into one reward function, enabling optimization of various and conflicting objectives during training (Okano et al., 2023; Pérez-Rosas et al., 2024; Dann et al., 2023). However, these approaches face significant challenges: maintaining training efficiency, ensuring scalability as objectives increase, and optimizing appropriate weights for each objective while preserving result explainability (Hayes et al., 2022; Dulac-Arnold et al., 2021). These challenges highlight the need for more efficient and flexible fine-tuning methods.
@@
Figure 1: The counselor reflection generation task requires creating reflective, empathetic, and fluent statements in response to the seeker’s problem.
Figure 2: Comparison of training efficiency (x-axis) and mean reward (y-axis). Point size and "+/-" indicate scalability and explainability capabilities. EMORL achieves comparable mean rewards with lower data consumption while maintaining better scalability and explainability.
@@
Ensemble learning offers a potential solution for this growing need by aggregating multiple trained models into a global model (Vanhaesebrouck et al., 2017). We propose a novel Ensemble Multi-Objective RL (EMORL) framework for LLM fine-tuning that distributes objectives to respective models, enabling training for individual objectives independently. The results demonstrate that models trained with single-objective reward functions converge significantly faster than those optimizing multiple objectives simultaneously. The framework incorporates a hidden states aggregation method coupled with a hierarchical grid search algorithm during aggregation, which efficiently identifies optimal weights for combining these single-objective models. Our experiments demonstrate that the aggregated output achieves higher training efficiency while achieving performance comparable to models trained using single-policy methods, as shown in Figure 2. This framework is also highly scalable, allowing additional objectives to be incorporated as modular components. Additionally, the framework enhances explainability by providing insights into the relative importance of different objectives easily through evaluating different weighted combinations of objectives on test samples. The code for our framework and experiments is publicly available and can be found at https://github.com/engineerkong/EMORL.
@@
Succinctly, our main contributions are as follows: (1) We introduce an ensemble multi-objective RL (EMORL) framework that separately trains and aggregates models in the counselor reflection generation task. (2) We develop an effective hidden-state level aggregation method and a hierarchical grid search algorithm for optimizing the weighted combination. (3) We demonstrate our framework’s effectiveness through a comprehensive evaluation against multi-objective baselines, achieving comparable performance across multiple objectives while offering lower and more stable training resources, improved scalability and explainability.
@@ 2 Related Work
Prior work on RL for LLM fine-tuning has explored various approaches to balance multiple objectives. The most common approaches use a single policy to learn all objectives simultaneously, requiring careful objective weight selection and balancing within the reward function or loss function. Below, we analyze different techniques and their respective advantages and disadvantages. Additionally, we introduce ensemble learning methods, which combine multiple policies to create a meta-policy that addresses multiple objectives.
@@ 2.1 Single-policy
In the most conventional approaches, objectives are combined into one reward function or loss function using fixed weights $\lambda_{i}$ for individual objectives, as shown in Equation (1), where the total reward $R_{total}$ is summed up using respective rewards $R_{i}$.
@@
Ziegler et al. (2019) explored RL for LLM fine-tuning using fixed weights to combine task-specific rewards with auxiliary objectives like fluency. However, selecting appropriate weights to balance objectives effectively is challenging, since it requires extensive trial-and-error to verify the effectiveness of weighted combinations. Mohan et al. (2023) proposed AutoRL to automate the selection of optimal weights. However, it still requires training numerous candidate policies to trace back the optima, remaining computationally intensive.
Dynamic weighting methods adjust objective weights during training based on the model’s performance, context, or external feedback, enabling a more flexible and adaptive balance between competing goals. Reward-driven approaches like Pérez-Rosas et al. (2024)’s Multi-Armed Bandit and Pu et al. (2024)’s Markov Chain strategies enable continuous adaptation based on received rewards. However, this approach requires these additional mechanisms to adjust weights, which increases computational complexity and introduces training instability. These limitations are evident in Figure 3, where the dynamic weighting method DynaOpt exhibits higher training costs and greater instability than the Uniform Weighted approach.
@@ 2.2 Meta-policy
To address these inefficiencies, meta-policy approaches use ensemble learning to reduce computational overhead and complexity. Ensemble learning comprises three main approaches: bagging (combining separately trained models), boosting (sequential training to improve upon previous models), and stacking (using a meta-learner to integrate outputs from diverse models, requiring additional meta-training) (Song et al., 2023). Wortsman et al. (2022) demonstrated Model Soups that combines the parameters from multiple trained models using the concept of bagging. Their learned soup approach learns the weights for parameter-level aggregation and aggregates multiple models into a single model through gradient-based optimization. Shi et al. (2024) combined the predicted tokens from models trained on individual objectives, namely logit-level aggregation, and leveraged Legendre transformation and f-divergence to optimize the performance of combined predictions for NLP tasks. Feriani et al. (2022) balanced the minimum throughput and standard deviation in load balancing by learning a meta-policy using distilled data from multiple trained models. Malik et al. (2024) trained a meta-policy by stacking three types of feature embedding representations from different trained models and used a multi-head attention mechanism to non-linearly combine the features for scoring sentiment analysis.
These meta-policy approaches shift the objective balancing from the training phase to the model aggregation, making the multi-objective optimization more flexible. However, as demonstrated in Appendix A.1 and  A.2, parameter-level aggregation and logit-level aggregation underperform in multi-objective optimization, particularly when the objectives differ significantly. This indicates that the application of ensemble learning in this multi-objective context remains underexplored. In light of this, we explore the novel hidden-state level aggregation by combining the contextual information from individual objectives to achieve efficient and flexible multi-objective LLM fine-tuning, while maintaining comparable performance.
@@ 3 Challenges
We analyze the challenges of multi-objective optimization in this section by comparing the training processes of models with individual objectives against multi-objective models using conventional single-policy methods. We tested five fine-tuning setups for counselor reflection generation to demonstrate these challenges, focusing on the objectives: reflection, empathy, and fluency. Three setups optimized single objectives separately, while two multi-objective approaches fine-tuned for all three objectives: (1) Uniform Weighted, assigning equal weights ($\frac{1}{3}$) to each objective, and (2) DynaOpt, dynamically adjusting weights using a multi-armed bandit algorithm. Experiments used $5$ random seeds with $3$ generation runs each, evaluating mean rewards, data consumption, and training time. Progress was tracked via Weights & Biases, plotting mean reward against data consumption.
As shown in Figure 3 and Table 3, our results highlight key differences between single-objective and conventional multi-objective fine-tuning in three aspects. First, convergence speed: single-objective models converged faster, with fluency models achieving the quickest convergence ($4,809$ data points, $1,629.19$ seconds). Multi-objective models were significantly slower, with Uniform Weighted requiring $23,398$ data points and $5,967.84$ seconds, and DynaOpt needing $31,328$ data points and $8,029.15$ seconds due to the overhead of dynamic weighting. Second, process stability: single-objective fine-tuning showed consistent convergence with minimal variation (±$\pm 335$ data points, ±$\pm 104.40$ seconds for reflection). Multi-objective models exhibited less stability, with Uniform Weighted varying by ±$\pm 7,390$ data points and ±$\pm 1,875.50$ seconds, and DynaOpt by ±$\pm 16,736$ data points and ±$\pm 4,365.64$ seconds, reflecting the complexity of balancing multiple objectives during training. Third, performance metrics: single-objective models achieved higher rewards (approaching $1.0$ for reflection and empathy), while multi-objective models averaged below $0.85$, indicating inherent performance trade-offs in optimizing multiple objectives simultaneously.
@@
These observations suggest that integrating multiple objectives inherently presents convergence, stability, and performance challenges. A promising research direction emerges from this insight: ensembling single-objective models to achieve multi-objective optimization.
@@ 4 Methodology
Figure 4: The EMORL framework illustrates a three-stage process: training, aggregation and inference.
As shown in Figure 4, the EMORL framework consists of three key stages. First, multiple models are independently fine-tuned for distinct individual objectives during training. Second, we employ a simple yet effective hierarchical grid search to explore optimal linear weighted combinations for aggregating these trained models at the hidden-state level. Finally, we use the optimal weights to aggregate the models at the inference stage, generating outputs that effectively integrate all objectives. This ensemble learning approach simplifies the complex multi-objective fine-tuning problem into an optimization problem for aggregation weights.
@@ 4.1 Hidden-State Level Aggregation
In contrast to parameter-level and logit-level aggregation methods, whose experimental results we present in Appendix A.1 and A.2 respectively, our hidden-state level aggregation approach ensures more consistent text generation while effectively incorporating features from all objective-specific models. The weight coefficients in this linear combination determine each objective’s relative contribution of contextual information to the final output.
@@ 4.2 Hierarchical Grid Search
We propose a hierarchical grid search algorithm that incorporates binary search concepts to reduce the computational complexity inherent in standard grid search (Bishop and Nasrabadi, 2006; Gautschi, 2011). The developed hierarchical grid search achieves a computational complexity of $O(3^{d}\cdot\log_{2}\frac{1}{N})$ compared to grid search’s $O(\frac{1}{N}^{d})$, where $d$ represents the objectives and $N$ the precision level. The complexity comparison among these methods is shown in Figure 9. With three objectives and a precision level of $0.03125$, the required number of evaluations is reduced to $135$, compared to $32,768$ in standard grid search.
As detailed in Algorithm 1, our hierarchical approach first divides each search axis for individual objectives into $3$ parts, creating $3^{d}$ initial grid points. We then evaluate the generation performance at these points and identify the most promising region by finding the $2^{d}$ cube with the highest total performance score. This region becomes the next search space, and we iterate these processes of grid generation and space refinement. The algorithm progressively focuses on smaller, more promising regions, which proves particularly effective for this aggregation case.
@@
1:objective function $f$, number of components $N$, iterations $I$, initial bounds $B_{0}=[(0,1)]^{N}$

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) for large language model (LLM) fine-tuning show promise in addressing multi-objective tasks but still face significant challenges, including competing objective balancing, low training efficiency, poor scalability, and limited explainability. Leveraging ensemble learning principles, we introduce an Ensemble Multi-Objective RL (EMORL) framework that fine-tunes multiple models with individual objectives while optimizing their aggregation after the fine-tuning to improve efficiency and flexibility. Our method is the first to aggregate the hidden states of individual models, incorporating contextual information from multiple objectives. This approach is supported by a hierarchical grid search algorithm that identifies optimal weighted combinations. We evaluate EMORL on counselor reflection generation tasks, using text classification models to score the generations and provide rewards during RL fine-tuning. Through comprehensive experiments on the PAIR and Psych8k datasets, we demonstrate the advantages of EMORL against existing baselines: significantly lower and more stable training consumption ($17,529\pm 1,650$ data points and $6,573\pm 147.43$ seconds), improved scalability and explainability, and comparable performance across multiple objectives.
@@
Lingxiao Kong1,*, Cong Yang2, Susanne Neufang3, Oya Deniz Beyan1,3, Zeyd Boukhers1,3,* 1Fraunhofer Institute for Applied Information Technology FIT, Germany 2Soochow University, China 3University Hospital of Cologne, Germany {lingxiao.kong, oya.deniz.beyan, zeyd.boukhers}@fit.fraunhofer.de susanne.neufang@uk-koeln.de cong.yang@suda.edu.cn
footnotetext: * Corresponding authors
@@ 1 Introduction
Figure 1: The counselor reflection generation task requires creating reflective, empathetic, and fluent statements in response to the patient’s problem.
Multi-objective optimization (MOO) for large language models (LLMs) is a crucial research direction in natural language processing (NLP) tasks that require fulfilling diverse and often competing requirements (Vaswani et al., 2017; Qin et al., 2024; Kumar, 2024). Such tasks require models to simultaneously optimize for multiple evaluation criteria, balancing trade-offs between different objectives. This paper focuses on the counselor reflection generation task, where LLM reflects on users’ prompts, building the first step of therapeutic communication (O’neil et al., 2023). As exemplified in Figure 1, this task requires generations that optimize competing objectives, such as reflection, empathy, and fluency in high-quality counseling interactions.
Reinforcement learning (RL) offers a promising approach for fine-tuning pre-trained LLMs in MOO (Lin, 2024; Parthasarathy et al., 2024). Conventional RL fine-tuning approaches combine multiple objectives into one reward function, enabling optimization of objectives during training (Okano et al., 2023; Min et al., 2024; Dann et al., 2023). However, these approaches face significant challenges: optimizing appropriate weights for each objective, maintaining training efficiency, and ensuring scalability while preserving result explainability (Hayes et al., 2022; Dulac-Arnold et al., 2021). From another perspective, compared to RL with Human Feedback (RLHF) (Ouyang et al., 2022), MOO explicitly defines objectives, advantaging user-specified LLM characteristics. Human feedback typically expresses preferences rather than specific objective values, requiring exhaustive human annotations, while MOO directly targets explicit objectives, transcending implicit instruction limitations. Challenges also exist in ensuring objectives align with human perception and in incorporating many objectives simultaneously. These challenges highlight the need for more efficient and flexible fine-tuning methods.
@@
Ensemble learning offers a potential solution for this growing need by aggregating multiple trained models into a global model (Vanhaesebrouck et al., 2017). We propose a novel Ensemble Multi-Objective RL (EMORL) framework for LLM fine-tuning that distributes objectives to respective models, enabling training for individual objectives independently. Our experiments demonstrate that models trained with single-objective reward functions converge significantly faster than those optimizing multiple objectives simultaneously. The framework incorporates a hidden states aggregation method coupled with a hierarchical grid search algorithm during aggregation, which efficiently identifies optimal weights for combining these single-objective models. Our results demonstrate that the aggregated output achieves higher training efficiency while achieving performance comparable to models trained using single-policy methods, as shown in Figure 2. The performance is represented by an objective reward, which uses a utility function that weights objectives according to user-specific preferences. Our framework is highly scalable, accommodating different user preferences and new objectives through aggregation. Additionally, the framework enhances explainability by providing insights into the importance of different objectives easily through evaluating different weighted combinations of objectives on test samples. The code for our framework and experiments is publicly available and can be found at https://github.com/engineerkong/EMORL.
@@
Succinctly, our main contributions are as follows: (1) We introduce EMORL framework that separately trains and aggregates models in the counselor reflection generation task. (2) We develop an effective hidden-state level aggregation method and a hierarchical grid search algorithm for optimizing the weighted combination. (3) We demonstrate our framework’s effectiveness through a comprehensive evaluation against MOO baselines, achieving comparable performance across multiple objectives while offering lower and more stable training resources, improved scalability and explainability.
Figure 2: Comparison of training consumption (x-axis) and objective reward (y-axis). Point size and "+/-" indicate scalability and explainability capabilities. EMORL achieves comparable performance with higher efficiency while maintaining better scalability and explainability.
@@ 2 Related Work
Prior work on RL for LLM fine-tuning has explored various approaches to balance multiple objectives. The most common methods include single-policy approaches that learn all objectives simultaneously using scalarized reward functions, and multi-policy approaches that train and populate numerous policies with vectorized rewards to achieve Pareto optimality. To address the efficiency and flexibility limitations of these existing methods, we analyze related single-policy and multi-policy techniques and present meta-policy methods with ensemble learning, which combines multiple policies to create a meta-policy that better addresses MOO.
@@ 2.1 Single-Policy
With prior knowledge, objectives are combined into one reward function or loss function using fixed weights $\lambda_{i}$ for individual objectives, as shown in Equation (1), where the total reward $R_{total}$ is summed up using respective rewards $R_{i}$.
@@
Ziegler et al. (2019) explored RL for LLM fine-tuning using fixed weights to combine task-specific rewards with auxiliary objectives like fluency. However, selecting appropriate weights to balance objectives effectively is challenging, since it requires extensive trial-and-error to verify the effectiveness of weighted combinations. Mohan et al. (2023) proposed AutoRL to automate the selection of optimal weights. However, it still requires training numerous candidate policies to trace back the optima, remaining computationally intensive. Dynamic weighting approaches like Min et al. (2024)’s Multi-Armed Bandit and Pu et al. (2024)’s Markov Chain strategies enable continuous adaptation based on received rewards. These methods adjust objective weights during training based on the model’s performance, context, or external feedback, enabling a more flexible and adaptive balance between competing goals. However, this approach requires these additional mechanisms to adjust weights, which increases computational complexity and introduces training instability.
2.2 Multi-Policy
Multi-policy methods train multiple models simultaneously and generate Pareto-optimal solution sets using evolutionary algorithms or other population-based mechanisms (Vamplew et al., 2022). These methods utilize vectorized rewards as shown in Equation (2) rather than scalarized reward functions. Since vectorized rewards encode no user preferences, the methods first iteratively obtain diverse Pareto-optimal solutions, then apply a user preference to select a single optimal solution.
$\mathbf{R_{total}}=\begin{bmatrix}R_{1}&R_{2}&\cdots&R_{n}\end{bmatrix}^{\top}$ (2)
These approaches offer a more comprehensive exploration of the solution space compared to single-policy methods. However, they require training numerous models and computationally expensive population mechanisms, making them particularly challenging for large-scale models like LLMs (Hayes et al., 2022).
@@ 2.3 Meta-Policy
To address these inflexibilities and inefficiencies, meta-policy approaches use ensemble learning to enhance model adaptability and reduce computational overhead (Yuan and Lai, 2024). Ensemble learning comprises three main approaches: bagging (combining separately trained models), boosting (sequential training to improve upon previous models), and stacking (using a meta-learner to integrate outputs from diverse models, requiring additional meta-training) (Song et al., 2024). Meta-policy approaches divide MOO into two hierarchical phases (Zhang et al., 2023). At the lower level, several models are trained according to different sampled user preferences. At the higher level, these models are combined to address specific user preferences and produce a single optimal solution. These approaches shift objective balancing from the lower level to the higher level, making multi-objective optimization more flexible and efficient.
Wortsman et al. (2022) demonstrated Model Soups that combines the parameters from multiple trained models using bagging. Their learned soup approach learns the weights for parameter-level aggregation and aggregates multiple models into a single model through gradient-based optimization. Shi et al. (2024) combined the predicted tokens from models trained on individual objectives, namely logit-level aggregation, and leveraged Legendre transformation and f-divergence to optimize the performance of combined predictions for NLP tasks. Feriani et al. (2022) balanced the minimum throughput and standard deviation in load balancing by learning a meta-policy using distilled data from multiple trained models. Malik et al. (2024) trained a meta-policy by stacking three types of feature embedding representations from different trained models and used a multi-head attention mechanism to non-linearly combine the features for scoring sentiment analysis.
However, as demonstrated in Appendix A.1 and  A.2, parameter-level aggregation and logit-level aggregation underperform in MOO, particularly when the objectives differ significantly. This indicates that the application of ensemble learning in this multi-objective context remains underexplored. In light of this, we explore the novel hidden-state level aggregation by combining the contextual information from individual objectives to achieve efficient and flexible multi-objective LLM fine-tuning, while maintaining comparable performance.
@@ 3 Challenges
In this section, we analyze the challenges of MOO by comparing the training processes of models with individual objectives against multi-objective models using conventional single-policy methods. We tested five fine-tuning setups for counselor reflection generation to demonstrate these challenges, focusing on the objectives: reflection, empathy, and fluency. Three setups optimized single objectives separately, while two single-policy approaches fine-tuned for all three objectives: (1) Uniform Weighted, assigning equal weights ($\frac{1}{3}$) to each objective, and (2) DynaOpt, dynamically adjusting weights using a multi-armed bandit algorithm. Experiments used $5$ random seeds with $3$ generation runs each, evaluating objective rewards, data consumption, and training time. Progress was tracked via Weights & Biases, plotting objective reward against data consumption.
As shown in Figure 3 and the detailed results in Table 3 (Section A), our results highlight key differences between single-objective and conventional MOO in three aspects. First, convergence speed: single-objective models converged faster, with fluency models achieving the quickest convergence ($4,809$ data points, $1,629.19$ seconds). Multi-objective models were significantly slower, with Uniform Weighted requiring $23,398$ data points and $5,967.84$ seconds, and DynaOpt needing $31,328$ data points and $8,029.15$ seconds due to the overhead of dynamic weighting. Second, process stability: single-objective fine-tuning showed consistent convergence with minimal variation (±$\pm 335$ data points, ±$\pm 104.40$ seconds for reflection). Multi-objective models exhibited less stability, with Uniform Weighted varying by ±$\pm 7,390$ data points and ±$\pm 1,875.50$ seconds, and DynaOpt by ±$\pm 16,736$ data points and ±$\pm 4,365.64$ seconds, reflecting the complexity of balancing multiple objectives during training. Third, performance metrics: single-objective models achieved higher objective rewards (approaching $1.0$ for reflection and empathy), while multi-objective models averaged below $0.85$, indicating inherent performance trade-offs in optimizing multiple objectives.
@@
These observations suggest that integrating multiple objectives in training inherently presents convergence, stability, and performance challenges. A promising research direction emerges from this insight: ensembling single-objective models to achieve efficient and flexible MOO.
@@ 4 Methodology
Figure 4: The EMORL framework illustrates a three-stage process: fine-tuning, aggregation and inference.
As shown in Figure 4, the EMORL framework consists of three key stages. First, multiple models are independently fine-tuned for distinct individual objectives. Second, we employ hierarchical grid search to find optimal linear combinations according to our utility function, which defaults to average objective preferences. Finally, we use the optimal weights to aggregate models at the hidden-state level during inference, generating outputs that effectively integrate all objectives. This ensemble learning approach simplifies the complex multi-objective fine-tuning problem into an optimization problem for aggregation weights.
@@ 4.1 Hidden-State Level Aggregation
In contrast to parameter-level and logit-level aggregation methods, whose experimental results we present in Appendix A.1 and A.2 respectively, our hidden-state level aggregation approach ensures more consistent text generation while effectively incorporating features from all objective-specific models. The weight coefficients in this linear combination determine each objective’s contribution of contextual information to the final output.
@@ 4.2 Hierarchical Grid Search
We propose a hierarchical grid search algorithm that incorporates binary search concepts to reduce the computational complexity inherent in standard grid search (Bishop and Nasrabadi, 2007; Gautschi, 2010). The developed hierarchical grid search achieves a computational complexity of $O(3^{d}\cdot\log_{2}\frac{1}{N})$ compared to grid search’s $O(\frac{1}{N}^{d})$, where $d$ represents the objectives and $N$ the precision level. The complexity comparison among these methods is shown in Figure 9. With three objectives and a precision level of $0.03125$, the required number of evaluations is reduced to $135$, compared to $32,768$ in standard grid search.
As detailed in Algorithm 1, we first define the utility function addressing our preferences for objectives. By default, we set the utility to average the three objective scores. Our hierarchical approach divides the search axes for individual objectives into $3$ parts, creating $3^{d}$ initial grid points. We evaluate generation performance at these points and identify the most promising region by finding the $2^{d}$ cube with the highest total performance that maximizes our utility function. This region becomes the next search space, and we iterate these processes of grid generation and space refinement. The algorithm progressively focuses on smaller, more promising regions, proving particularly effective for this aggregation case.
@@
1:utility function $f$, number of components $N$, iterations $I$, initial bounds $B_{0}=[(0,1)]^{N}$